OpenStudy (anonymous):

What would the equation of a line in standard form be if the equations were (1,2) and (2,3)?

OpenStudy (anonymous):

@Preetha ??

OpenStudy (anonymous):

1x - y = 1?

OpenStudy (debbieg):

Check your signs. You have a sign wrong... otherwise correct.

OpenStudy (debbieg):

(and you don't need to put the coefficient "1" in front of the x)

OpenStudy (anonymous):

Which sign do I have wrong? Can you explain why its wrong? @DebbieG

OpenStudy (debbieg):

Can you show your work? You started with the 2 points, (1,2) and (2,3).... what did you do first?

OpenStudy (anonymous):

y - (3) = 1(x - 2) y - 3 = 1x - 2 y - 3 = 1x - 2 -1x -1x -1x + y - 3 = -2 +3 +3 -1x + y = 1 1x - y = 1

OpenStudy (anonymous):

@DebbieG

OpenStudy (debbieg):

Good. it all looks good.... until the final step. What did you do, to get that pesky negative coefficient on the x, to be positive?

OpenStudy (anonymous):

I multiplied everything by -1

OpenStudy (debbieg):

Well, almost everything. Everything on the left side, you did. ;)

OpenStudy (anonymous):

-1! Thank you:)

OpenStudy (debbieg):

exactly! :) and you're welcome. :)

OpenStudy (anonymous):

@DebbieG Can I have your help with another question that i am a little confused about?

OpenStudy (debbieg):

I can try to help, but you should really post each question separately. It looks like you're offline now and I have to go do some other things, but just post it as a new question, and someone will help! :) You can always tag me if you want, and if I'm online I'll try to help.